# Feature Identification via the Empirical NTK

### v1

@@ Abstract
We provide evidence that eigenanalysis of the empirical neural tangent kernel (eNTK) can surface the features used by trained neural networks. Across two standard toy models for mechanistic interpretability, Toy Models of Superposition (TMS) and a 1-layer MLP trained on modular addition, we find that the eNTK exhibits sharp spectral cliffs whose top eigenspaces align with ground-truth features. In TMS, the eNTK recovers the ground-truth features in both the sparse (high superposition) and dense regimes. In modular arithmetic, the eNTK can be used to recover Fourier feature families. Moreover, we provide evidence that a layerwise eNTK localizes features to specific layers and that the evolution of the eNTK eigenspectrum can be used to diagnose the grokking phase transition. These results suggest that eNTK analysis may provide a practical handle for feature discovery and for detecting phase changes in small models.
@@ 1 Introduction
where the index $i$ runs over output neurons (classes), $\alpha$ runs over points in the training set $\mathcal{D}$, and the data-space kernel
@@
Related work. A few earlier works have explored whether NTK eigenvectors align with interesting directions in trained models. Loo et al. (2022) and Tsilivis and Kempe (2022) visualized kernel eigenvectors as candidate features in small image nets (in the context of using the NTK to study adversarial training), and are the most closely related works. Atanasov et al. (2021) and Baratin et al. (2021) argued that NTK tangent features rotate towards directions predictive of the outputs during training.
Besides the eNTK, other objects involving gradients/curvature have appeared in the mechanistic interpretability literature. As mentioned above, the eNTK (3) is constructed by contracting a pair of Jacobians $J(z)=\nabla_{W}f(z)$ along the parameter direction, forming a kernel in data space (shorthand: $K(z,z^{\prime})=J(z)J(z^{\prime})^{T}=\nabla_{W}f(z)\nabla_{W}f(z^{\prime})^{T}$). The Hessian, the second derivative of the loss function wrt parameters $H=\nabla^{2}_{W}\mathcal{L}$, can be approximated by a pair of Jacobians contracted the other way ($H\approx\sum_{i}J(x_{i})^{T}\,\nabla^{2}_{f}\mathcal{L}(f(x_{i}),y_{i})\,J(x_{i})$), and appears frequently in recent work on singular learning theory (e.g Bushnaq et al. (2024)). The Hessian induces a ``influence function" kernel on data points $\mathcal{I}(z^{\prime},z)=-\nabla_{W}\mathcal{L}(z^{\prime})^{T}H^{-1}\nabla_{W}\mathcal{L}(z)$ (Koh and Liang (2017)) that was recently used to interpret LLMs at scale (Grosse et al. (2023)). Somewhat further afield, attribution patching (Nanda (2023)) approximates activation patching (an interpretability technique where one swaps clean activations into a corrupted forward pass to test causal effect) by estimating the effect using the activation's local gradient.”
@@ 3.1 Review of the setup
The claim to fame of this setup is that it's a maximally simple toy model of superposition: i.e., the idea that if models store features as linear combinations of within-layer neurons, then a model might learn to represent more features than neurons available to it by storing them non-orthogonally in the neural activations. In this experiment, the $n$ input/output directions are the analogs of ground-truth features, and one would like to see if the autoencoder is able to reconstruct them. Upon training this class of model, one finds that when the features are dense (small $S$), the model learns to reconstruct the top $m$ features only, but as one dials up the sparsity, the model is able to recover more and more features, at the cost of encoding them non-orthogonally.
@@
Figure 1: Eigenvalue spectrum of the flattened eNTK (top row) and column-normalized, importance-rescaled flattened eNTK eigenvector-feature indicator heatmaps for TMS models with $n=50$ ground-truth features, importance $I_{i}=0.8^{i}$, (hidden-layer size, sparsity) $m=10,S=0.3$ (left column); $m=10,S=0.9$ (middle column) or $m=40,S=0.9$ (right column); and importance rescaling parameters $\beta=1$ (middle row) and $\beta=0.3$ (bottom row). The center column is an example of a TMS model in the sparse regime (high superposition), while the left and right columns are two different ways to push the model to the dense regime (low superposition).
@@ 3.2 Results
Another modification that we do ahead of rendering the heatmaps is to apply an importance rescaling to the eNTK, multiplying each Jacobian entry $df_{i}(x)/dW_{\mu}$ in (3) by $I^{\beta i/2}$. This directionally corrects for the fact that the derivation of Eq. (1) assumed a standard MSE loss (see Appendix A), but TMS uses an importance-weighted MSE loss, (5). The choice $\beta=1$ is the theoretically principled one that absorbs $I$ into the NTK in eq. (13), allowing the rest of the derivation in Appendix A to proceed with standard MSE, while $\beta=0$ corresponds to turning off the importance rescaling. With $\beta=1$, we find that the leading eigenvectors align with high-importance features (middle row of Fig. 1).
We also sweep over other values of $\beta\in[0,1]$ and find that as $\beta$ decreases, lower-importance features become increasingly visible in the heatmaps. In particular, at the value $\beta=0.3$ (bottom row of Fig. 1), the leading eigendirections of the importance-rescaled eNTK appear to align nearly one-to-one with all of the ground-truth features of TMS in both the sparse and dense regimes. This experiment suggests that the raw kernel also contains information about less-important features, but down-weighs them in line with the objective.
@@ 4.1 Review of the setup
After applying trigonometric identities, the second-layer activations turn out to be the sum of many terms each with the schematic form $\sum_{k=1}^{n}\cos(2\pi\frac{k}{p}s+\varphi_{k})$, for $s$ a linear combination of $\{a,b,q\}$ and $\varphi_{k}$ a linear combination of $\{\varphi_{k}^{(1)},\varphi_{k}^{(2)},\varphi_{k}^{(3)}\}$. For generic phases these sums cancel, but one such term turns out to use exactly the combination $\varphi_{k}^{(1)}+\varphi_{k}^{(2)}-\varphi_{k}^{(3)}=0$. Hence, its phase drops out and the remaining $\sum_{k=1}^{n}\cos(2\pi\frac{k}{p}s)$ acts as a modular $\delta$ function, firing exactly when s=a+bqs=a+b-q.
@@ 4.2 Results
On the other hand, the first cliff is present already at initialization (see the right panel of Fig. 3). This suggests (in combination with the subsection below) that this model's use of Fourier modes could perhaps have been predicted at initialization from the eNTK.
@@ Acknowledgements
I am grateful to Ari Brill, Lauren Greenspan, Andrew Mack, Logan Smith, and Dmitry Vaintrob for discussions and to Lauren Greenspan for comments on the manuscript. This work was supported by PIBBSS and the Long-Term Future Fund.
@@ Appendix B Kernels used in this paper
The total kernel is their sum, $K=K^{(1)}+K^{(2)}$. It could be interesting, and probably is tractable, to understand what properties of the weights and data lead to the mechanistic structure found in the main text. However, I'll leave this for future work.

### v2

@@ Abstract
We provide evidence that eigenanalysis of the empirical neural tangent kernel (eNTK) can surface the features used by trained neural networks. Across two standard toy models for mechanistic interpretability, Toy Models of Superposition (TMS) and a 1-layer MLP trained on modular addition, we find that the eNTK exhibits sharp spectral cliffs whose top eigenspaces align with ground-truth features. In TMS, the eNTK recovers the ground-truth features in both the sparse (high superposition) and dense regimes. In modular arithmetic, the eNTK can be used to recover Fourier feature families. Moreover, we provide evidence that a layerwise eNTK localizes features to specific layers and that the evolution of the eNTK spectrum can be used to diagnose the grokking phase transition. These results suggest that eNTK analysis may provide a practical handle for feature discovery and for detecting phase changes in small models.
@@ 1 Introduction
where the index $i$ runs over output neurons (classes), the $\alpha'$s run over points in the training set $\mathcal{D}$, and the data-space kernel
@@
Related work. A few earlier works explored whether NTK eigenvectors align with interesting directions in trained models. Loo et al. (2022) and Tsilivis and Kempe (2022) visualized kernel eigenvectors as candidate features in small image nets (in the context of using the NTK to study adversarial training), and are the most closely related works. Atanasov et al. (2021) and Baratin et al. (2021) argued that NTK tangent features rotate towards directions predictive of the outputs during training.
Besides the eNTK, other objects involving gradients/curvature have appeared in the mechanistic interpretability literature. As mentioned above, the eNTK (3) is constructed by contracting a pair of Jacobians $J(z)=\nabla_{W}f(z)$ along the parameter direction, forming a kernel in data space (shorthand: $K(z,z^{\prime})=J(z)J(z^{\prime})^{T}=\nabla_{W}f(z)\nabla_{W}f(z^{\prime})^{T}$). The Hessian, the second derivative of the loss function wrt parameters $H=\nabla^{2}_{W}\mathcal{L}$, can be approximated by a pair of Jacobians contracted the other way ($H\approx\sum_{i}J(x_{i})^{T}\,\nabla^{2}_{f}\mathcal{L}(f(x_{i}),y_{i})\,J(x_{i})$), and appears frequently in recent work on singular learning theory (e.g Bushnaq et al. (2024b)). The Hessian induces a ``influence function" kernel on data points $\mathcal{I}(z^{\prime},z)=-\nabla_{W}\mathcal{L}(z^{\prime})^{T}H^{-1}\nabla_{W}\mathcal{L}(z)$ (Koh and Liang (2017)) that was recently used to interpret LLMs at scale (Grosse et al. (2023)). The Jacobian also played a role in the ``Local Interaction Basis" algorithm (Bushnaq et al. (2024a)) which proposed to find interpretable directions in activation space in part by aligning would-be directions with singular directions of the intra-layer Jacobian. Somewhat further afield, attribution patching (Nanda (2023)) approximates activation patching (an interpretability technique where one swaps clean activations into a corrupted forward pass to test causal effect) by estimating the effect using the activation's local gradient.
Non-gradient-based kernels have also appeared in the interpretability literature. In SAE work (e.g. Bricken et al. (2023), Templeton et al. (2024)) it's common to compute a kernel-like measure of similarity between SAE features (e.g. cosine similarity of their neuron supports), then use nearest-neighbors/UMAP to visualize the feature manifold. More recently, motivated by singular learning theory, Adam et al. (2025) used a data-space ``loss kernel" defined as the covariance of per-sample losses under low-loss perturbations of a model's weights to cluster inputs/map dataset geometry.
@@ 3.1 Review of the setup
The claim to fame of this setup is that it's a maximally simple toy model of superposition: namely, the idea that if models store features as linear combinations of within-layer neurons, then a model might learn to represent more features than neurons available to it by storing them non-orthogonally in the neural activations. In this experiment, the $n$ input/output directions are the analogs of ground-truth features, and one would like to see if the autoencoder is able to reconstruct them. Upon training this class of model, one finds that when the features are dense (small $S$), the model learns to reconstruct the top $m$ features only, but as one dials up the sparsity, the model is able to recover more and more features, at the cost of encoding them non-orthogonally.
@@
Figure 1: Eigenvalue spectrum of the flattened eNTK (top row) and column-normalized, importance-rescaled flattened eNTK eigenvector-feature indicator heatmaps for TMS models with $n=50$ ground-truth features, importance $I_{i}=0.8^{i}$, (hidden-layer size, sparsity) $m=10,S=0.3$ (left column); $m=10,S=0.9$ (middle column) or $m=40,S=0.9$ (right column); and importance rescaling parameters $\beta=1$ (middle row) and $\beta=0.3$ (bottom row) used as described in the main text. The center column is an example of a TMS model in the sparse regime (high superposition), while the left and right columns are two different ways to push the model to the dense regime (low superposition).
@@ 3.2 Results
Before rendering the heatmaps, we also apply an importance rescaling to the eNTK, multiplying each Jacobian entry $df_{i}(x)/dW_{\mu}$ in (3) by $I^{\beta i/2}$. This directionally corrects for the fact that the derivation of Eq. (1) assumed a standard MSE loss (see Appendix A), but TMS uses an importance-weighted MSE loss, (5). The choice $\beta=1$ is the theoretically principled one that absorbs $I$ into the NTK in eq. (13), allowing the rest of the derivation in Appendix A to proceed with standard MSE, while $\beta=0$ corresponds to turning off the importance rescaling.
With $\beta=1$, we find that the leading eigenvectors align with high-importance features (middle row of Fig. 1). We also sweep over other values of $\beta\in[0,1]$ and find that as $\beta$ decreases, lower-importance features become increasingly visible in the heatmaps. In particular, at the value $\beta=0.3$ (bottom row of Fig. 1), the leading eigendirections of the importance-rescaled eNTK appear to align nearly one-to-one with all of the ground-truth features of TMS in both the sparse and dense regimes. This experiment suggests that the raw kernel also contains information about less-important features, but down-weighs them in line with the objective.
@@ 4.1 Review of the setup
After applying trigonometric identities, the second-layer activations turn out to be the sum of many terms each with the schematic form $\sum_{k=1}^{n}\cos(2\pi\frac{k}{p}s+\varphi_{k})$, for $s$ a linear combination of $\{a,b,q\}$ and $\varphi_{k}$ a linear combination of $\{\varphi_{k}^{(1)},\varphi_{k}^{(2)},\varphi_{k}^{(3)}\}$. For generic phases these sums cancel, but one such term turns out to use exactly the combination $\varphi_{k}^{(1)}+\varphi_{k}^{(2)}-\varphi_{k}^{(3)}=0$. Hence, its phase drops out and the remaining $\sum_{k=1}^{n}\cos(2\pi\frac{k}{p}s)$ acts as a modular $\delta$ function, firing when s=a+bqs=a+b-q.
@@ 4.2 Results
On the other hand, the first cliff is present already at initialization (see the r.h.s. of Fig. 3). This follows from the structure of the Layer 1 kernel, as we explain briefly in Appendix B, and suggests (in combination with the subsection below) that this model's use of Fourier features could perhaps have been predicted at initialization from the eNTK.
@@ Acknowledgements
I am grateful to Ari Brill, Lauren Greenspan, Andrew Mack, Logan Smith, Lucas Teixeira and Dmitry Vaintrob for discussions and comments on the manuscript. This work was supported by PIBBSS and the Long-Term Future Fund.
@@ Appendix B Kernels used in this paper
The total kernel is their sum, $K=K^{(1)}+K^{(2)}$.
It could be interesting, and probably is tractable, to understand which properties of the weights and data lead to the mechanistic structure found in the main text. In general, I will leave this for future work.
As an appetizer however, note that an especially striking result in the main text - the fact that the Layer 1 eigenvalue cliff in the modular arithmetic problem was present already at initialization, and hence that the use of Fourier features for modular arithmetic could perhaps have been predicted at initialization with the eNTK - is easily explained by the structure of Eq. (18). Eq. (18) contains a parameter-agnostic factor of $x_{1}\cdot x_{2}$ which is zero for one-hot-encoded data points $(a,b)$ and $(a^{\prime},b^{\prime})$ unless either $a=a^{\prime}$ or $b=b^{\prime}$. This makes the Layer 1 kernel have nonzero values along each fixed-$a$ and fixed-$b$ line, and zeroes elsewhere. Moreover, the nonzero kernel entries at initialization are approximately identical for each choice of $a$ or $b$ by symmetry. Hence, the kernel at initialization has an approximate shift symmetry along the $a$ and $b$ axes, so is approximately diagonalized by Fourier waves on each axis, with (nonzero) eigenvalues related to the weight initializations.
In this example, the $x_{1}\cdot x_{2}$ factor in (18) relied on the model architecture (specifically, the fact that each one-hot-encoded data point only touches two columns of the Layer 1 weight matrix), and its conversion to a shift-symmetric kernel relied on the shift symmetry of the dataset, illustrating that the NTK (and its eigenstructure) inherits information about both already at initialization.